[Brahms-l] Shift report 20040206 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Fri Feb 06 2004 - 16:07:57 EST
------------------------------------------------------------
Date: Fri Feb  6 16:07:57 2004, Shift: 08:00-16:00
Supervisor: rk, with: 
Log book#15 pages 49 - 49

Shift Summary:
------------------------------------------------------------
1. Beam for around 4 hours, but due to various reasons only about 3 hours were used.
2. Only two data runs: 10183 and 10184 at FS@8@1723B and MRS@45@2255A.
FS was moved to 2.3 degrees.
3. NEW MAGNET SOFTWARE (see E-log for detail). Don't use the old magnet software.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 6 16:08:10 2004

This archive was generated by hypermail 2.1.8 : Fri Feb 06 2004 - 16:08:27 EST